Q.1] The perimeter of a sector of a circle of area 36pie sq.cms is 24cms. Find the area of the sector.


Q.2] The perimeter of a sector of a circle is 4 times the radius of the circle. find the angle of the sector in radians and degrees.


Q.3] A wire 121 cm long is bent so as to lie along the arc of a circle of 180 cm radius. Find in degrees,the angle subtended at the centre of the arc.


Q.4] A horse is tied to a pole by a rope 30 m long. If the horse moves along the circumference of a circle always keeping the rope tight, find how far it would move when the rope has traced an angle of 105 degree ?

Grade:Upto college level

1 Answers

Ramesh V
14 years ago

1) perimeter = 2.Pi.r

area = Pi.r2

so , r=6cms

so x.2.Pi.6 / 360 = 24

so sector angle = 229.3 deg

area of sectore = 229.3/360 * 36Pi  = 22.93.Pi

area of sectore = 22.93.Pi

 

2) x.2.Pi.r /360 = 4* r

x = 229.3 deg or 4 radians

3)

x.2.Pi.180 /360 = 121

angle subtended at centre: x = 38.22 deg

4)

105.2.Pi.30 /360 = 54.95 mts

--

Please feel free to post as many doubts on our disucssion forum as you can. If you find any question difficult to understand - post it here and we will get you the answer and detailed solution very quickly.We are all IITians and here to help you in your IIT JEE preparation. All the best.

Regards,
Naga Ramesh
IIT Kgp - 2005 batch